1 Introduction

This document describes the policies, guidelines, and traditions of French House. Although the idea of writing such a document has been proposed for several years, it was not until the spring of 1993 that the first draft was actually written. This is not a “Constitution” per se, although rules and policies are official and everyone is expected to abide by them.

All rules and policies remain in effect until removed or altered at a house meeting.

This document was last revised July 31, 2001.

2 Membership

2.1 Resident Members

Any person living in French House is a resident member of French House. The only exception to this rule is a student who was placed in French House by the MIT administration without prior consent of the house, and who has no intention of sharing the responsibilities which are required of resident members.

2.2 Social Members

A social member is defined as a person who does not live in French House, but who is interested in French House and who participates in the duties and activities of the house. Social membership in French House must be approved by the rest of the house members at a house meeting. Prospective social members must not be present during the discussion and vote on their membership. Once accepted, social members retain their house membership until they decide to relinquish it. In the extreme and unlikely case that a social member does not fulfill the requisite duties, the house may vote to revoke membership.

2.3 Graduate Resident Tutor

The tutor is a graduate student who lives in French House and who is thus considered a resident member of the house. The role of the tutor is somewhat vague and subject to individual opinion, but may include the following:

• To serve as a liaison between French House and the MIT administration

• To encourage the use of French in the house

• To suggest and/or organize French House activities

• To act as an impartial mediator in house or personal disputes

• To offer a sympathetic ear, and perhaps advice and counsel, to house members who need it

The tutor is selected by house vote at a special tutor meeting, after the house has had an opportunity to interview tutor candidates. Again, the basis of tutor selection is subjective and individual, but should have some basis in French ability and the attributes listed above. The house’s tutor selection is subject to final approval by the Dean’s Office.

3 French

Mais bien sûr, c’est La Maison Française. Nous parlons français.

French is to be spoken during dîners and pauses-café. The only time you are not required to speak French during dinner is when there is a guest at the table who does not understand French. Then all seated at that table may speak English.

French is also spoken during house meetings.

During French marathons, French is spoken continuously.

Members are encouraged to speak French at any other time.^{1}

4 Duties

4.1 Cooking

Every member of French House has several duties to fulfill, and perhaps the most important duty is cooking. Almost every member cooks on a cooking team once per week. The Ministre des Aliments and Trésorier are exempt from cooking because these positions require a great deal of time. Also, when the house membership is large the house may decide to have a dedicated Grand Nettoyage team which cleans once per week. (See Section 4.3.1 below). Please see Section 7.1.5 for more details.

If you cannot cook on a specific day, it is your responsibility to find someone to trade days with you or make alternate arrangements with your team.

4.2 Tâches

Resident members of French House are assigned tâches (chores) at the beginning of the semester. Tâches are an important part of keeping the house a little tidier or running smoothly and are not to be ignored. Most tâches are randomly assigned, but a few (bringing up the morning paper, bringing up the milk order, sorting the vegetable order, sorting the meat order, and washing the dishtowels) require a bit more effort or have specific time commitments. The Président asks for volunteers for these tâches.

4.3 Nettoyage

4.3.1 Grand Nettoyage

Grand Nettoyage involves cleaning refrigerators, counters, stoves, and pantries. It is done weekly, usually over a weekend. How and by whom is determined at the beginning of the semester.

4.3.2 Nettoyage Enorme

The Nettoyage Enorme happens at least twice a year and is organized by the Président. It takes place once during the days before Rush begins, and once during IAP. There may also be a Nettoyage Enorme at the end of the second semester, at the Président’s discretion. Everyone has a specific job assigned to them to give the house and the kitchen in particular a thorough cleaning.

1Occasionally people actually get annoyed if you speak French, but don’t let that discourage you. Speak away!

4.4 Recycling

House members are asked to recycle what they can. Tin cans, soda cans and bottles, etc., should be rinsed out well before being tossed in the recycling bins. The Ministre de l’Environnement is responsible for coordinating French House’s particpation in New House recycling. Emptying the bins, however, is one of the tâches.

4.5 Implementation Policies

The house may institute duty implementation policies to discourage delinquency in the payment of housebills and to encourage fulfillment of cooking and nettoyage duties. Such policies may require the Trésorier to assess, and Vice-Président to record, monetary and non-monetary penalties, respectively.

5 Etiquette

Here are a few unspoken “rules” of courtesy that French House members are strongly encouraged to follow.

• When you want something out of the newspaper (a comic strip, an article, etc.), don’t cut it out for a day. It is a good idea to mark on it that you want it!

• Water fights are to be kept out of people’s rooms and away from posters which may be damaged by water.

• If food is left out for the house to eat, the person who takes the last portion must clean up the container. Please don’t just divide the remaining portion in half in an attempt to avoid the simple task of cleaning; this results in a dirty, essentially empty container and several annoyed house members. The same rule applies to leftovers in the refrigerators.

• It is considerate for cooking team members not to be the first ones in line for dinner.

6 Kitchen

6.1 Using the Kitchen

The kitchen is open for use any time (although it is polite to avoid cooking while a cooking team is preparing a meal or cleaning). House kitchen equipment can be freely utilized, but must be cleaned after use. Also be sure to clean tables, stoves, and counters if you have dirtied them.

Don’t leave dirty pots or utensils sitting on the counter or in the sink. Please be especially careful not to leave silverware or other small utensils in the sink, lest they accidentally fall into the garbage disposal. If you are planning to return later to clean something, please be considerate and leave a note stating that you will return to clean it. Cleaning should be done as promptly as possible.

6.2 Dishwasher

Dishes should always be rinsed after use, and perhaps even scrubbed. They should then be put in the dishwasher if it is not running. If it has finished washing dishes and has not yet been emptied, please empty it. Then put dirty dishes that have accumulated by the sink into the dishwasher. If the dishwasher is full, fill both detergent cups with detergent and start the dishwasher.

There is a standard method for communicating whether the dishes in the dishwasher are clean or not. If the handle is in the locked position, it is running or has been run. If the handle is the locked position and it is not running, it is waiting for you to empty it. Leave the handle unlocked at all other times to avoid confusing others. No one wants to eat off of dirty plates!

Pots and pans should be washed by hand, not in the dishwasher. Do not put any sharp knives in the dishwasher. Finally, keep wooden utensils (especially knives with wooden handles!) out of the dishwasher.

6.3 Storage

6.3.1 Refrigerators

Two refrigerators are provided for storing personal food. It is advisable to label all personal food, since unlabeled food in the personal refrigerators may be eaten by anyone. Avoid storing food on house dishes.

Personal food is not to be stored in the two house refrigerators. (That is why they are called “house fridges” and not “personal fridges!”) Food in these may be eaten even if it is labeled with someone’s name.^{2}

6.3.2 Cupboards

In theory, everyone gets a shelf in a cupboard for storing personal food. If the house is large this may be a problem. Please compromise and share. Everything in cupboards should also be labeled.

7 Food

7.1 Dîner and Pause-Café

7.1.1 Dinner Schedule

French House traditionally serves dinner every night excluding Saturday. Dinner is served at a specified time which in the past has been somewhere between 6 and 6:30 PM; usually Friday and Sunday dinners are served a bit earlier than on other days. The actual schedule of dinner days and times is to be determined at the beginning of each semester.

There is no dinner served during long holidays (any holiday lasting longer than a normal weekend), except possibly on the day before classes resume. That day’s cooking team will decide at least two weeks in advance if the team can cook on that day, and should consult with the Ministre des Aliments before making a final decision. The captain should determine if there will be enough cooking team members (or substitutes) available to cook. If so, then the captain should post a survey asking house members if they want dinner on that day. A dinner will be prepared if a majority of the house wants to have dinner.

7.1.2 Pause-Café

Every week a member of the house prepares a dessert or snack food for a pause-café (study break). Study breaks are usually held in the middle of the week, around 9-10PM. The study break chef may determine the actual day, although it should be consistent throughout the term.

2The exceptions to this rule are gardés. See Section 7.1.3.

7.1.3 Dinner Sheets

A dinner sheet is posted by the Trésorier each week. It is a grid with cooking days and study breaks listed horizontally and house members listed vertically. Dinner sheets are used by cooking teams to determine how many people to expect for dinner, and by the Trésorier for bookkeeping purposes.

There are four letters to write on the dinner sheets: X, G, I, and V. X indicates that you did not eat dinner that day;^{3} G (gardé) indicates that you wish the cooking team to save a portion for you (a.k.a. late plate); I (invité) indicates that you plan to have a guest for dinner; and V (végétarien) means that you are eating a vegetarian meal (i.e. not eating the meat portion of the dinner) and should therefore be charged the vegetarian dinner price instead of the meat dinner price. A blank space indicates that you ate a normal meal.

7.1.4 Menus

Dinners and study breaks are described on menu sheets which are prepared by cooking team members and posted in the dining area weekly by the Ministre des Aliments. In addition to specifying the courses of the meal, menu sheets list the foods required for dinner preparation. Before posting menus, the Ministre des Aliments indicates on the sheets the foods that will be ordered and those which are already in the house. Any item not ordered and not available in the house must be bought by the cooking team (and usually by the person who wrote the menu).

When preparing a menu sheet, write down everything you expect to need, not just what you think the Ministre des Aliments can order. Please remember to prepare an appropriate amount of food for vegetarians and others with dietary restrictions. Always remember to check menus once they are posted for anything which must be bought. It is also a good idea to verify that foods listed as “in the house” and “ordered” are actually available in sufficient quantities.

Label receipts from food purchases with the menu date, the buyer’s name, and the total amount, and give them to the Trésorier. The money you spent will be credited to your house account. The Ministre des Aliments reserves the right to determine beforehand a reasonable limit to the amount of money spent on non-ordered food.

The due date for menus is decided by the Ministre des Aliments at the beginning of the semester. Please ask your cooking team captain or the Ministre des Aliments if you have any questions regarding menus.

7.1.5 Cooking

Cooking Teams The Ministre des Aliments will post a signup sheet at the beginning of each semester, and ask that everyone rank their top n choices for cooking days (including study breaks). When everyone has indicated their preferences for cooking day, the Ministre des Aliments will assign cooking teams. The main goal is to balance cooking team size (optimally four members). In addition, for the first semester each cooking team should have at least one upperclass member who has been on a cooking team before. After these considerations are taken into account the Ministre des Aliments will choose the cooking teams so as to give the most people their top choices.

3Please don’t think it rude to mark an X before dinner rather than after. Cooking teams will appreciate knowing that you did not plan to eat at the house.

Chefs The chef for a particular meal is a member of the cooking team who oversees the preparation of that meal and the cleanup afterward. The chef is almost always the person who has written the menu sheet for the dinner. Cooking teams normally distribute among themselves as equally as possible the number of menus each person must write and hence the number of times each team member is chef. However, occasionally a cooking team may decide (with the consent of the Ministre des Aliments) to collaborate on menu-writing and dispense with actual chefs.

7.1.6 Uncooked Leftovers

Rule of thumb for cooking teams: If you could not eat the uncooked leftovers by yourself in one sitting, they should be referred to the Ministre des Aliments. Otherwise, mark them “eat me.”

7.2 What Can I Eat?

7.2.1 Signout Food

As a convenience, French House orders several foods and makes them available to house members on a signout basis (such as milk, fruit juices, iced tea, and sometimes breads and pastries). When a house member wishes to use these foods, it is necessary to fill in the appropriate bubble on the signout sheets. The Trésorier then adjusts the house member’s account accordingly.

7.2.2 Non-Signout Food

Almost all other foods are for free general use (no signout needed), such as flour, pasta, oil, eggs, etc. Please use them in moderation (i.e. don’t feed 20 friends weekly). Such things as frozen bread, croissants, or bagels do NOT fall in this category (they must be signed out).

7.2.3 Food in the House Fridges

All food should be well labeled with a cooking day, “eat me,” “mangez-moi,” or that it’s waiting for a decision by the Ministre des Aliments. If food is unlabeled, it should be assumed to be waiting for reappropriation and labeling by the Ministre des Aliments, and so it should not be eaten. If it looks neglected or forgotten, it should be reported to the Ministre des Aliments. The food must be directly labeled. If you are unsure whether it is permissible to eat a certain food, please ask the Ministre des Aliments.

7.2.4 Dinner Food

If you have marked an X for a particular dinner, you may not eat food from that dinner until the following day (yes, this means wait until after midnight!). This is out of fairness to the people who have actually paid for the dinner.

8 IM Sports

French House participates actively in intramural sports. Usually we play in the lowest or next-to-lowest league. The goal is to have fun rather than to stress out about wins and losses.

Please don’t sign up for a sport if you are not going to be willing to take the time to play the games. Remember that by signing up for a sport you are committing the time to play.

8.1 IM Players

Anyone who signs up for a sport that the house plays is a player. Players are responsible for attending any games that they reasonably can. Players are not required to play just because they signed up, but if the team does not have enough players for a particular game, the house forfeits the game and is fined (usually $25 for the first forfeit, $50 for the second). The entire house pays these fines, not just the team members.

8.2 IM Captain

The Ministre des Sports will choose a captain for an IM team from a list of willing captains for that sport. The captain must usually attend one mandatory meeting before the season starts; the house is fined if the captain does not attend. The captain is responsible for rounding up players for games. The best way is to post signup sheets at least a week before the scheduled game, and to announce upcoming games at dinner. The captain has the same responsibilities as a regular player, and may be given other responsibilities due to the nature of the game.

8.3 IM Referee

For some IM sports, French House is required to provide referees. The Ministre des Sports chooses the referees from a list of volunteers. They must attend mandatory referee meetings; again, the house is fined if the referees miss the meeting. Referees must usually attend and referee three games per season, and missing these games results in fines. Referees have the same responsibilities as players, but should not be expected to play in as many games since they have additional responsibilities.

9 Housing Policy

Housing policy is one of the least understood and most often changed set of rules in the house. In recent memory there has been a reworking or attempted reworking of housing policy at least once every year. This has primarily been to deal with unexpected or unusual situations that arise.

In this housing policy section, for the purposes of discussion, “summer” is defined as the period of time between the beginning of the spring lottery and the end of the following Rush, and is not to be construed as referring to summer residents. “Non-summer” refers to all other times. For obvious reasons, the French House graduate tutor’s residence is not affected by the following policy.

NOTE: The Residence and Campus Activities Office (referred to as “the housing office”) reserves the right to overrule any housing actions or decisions.^{4}

9.1 Getting in to French House; Order of Preference

NOTE: This section deals only with the issues concerning who resides in French House. It does not pertain to the order in which rooms are chosen.

The following rules determine the order in which people are considered for residence in French House:

1. “Squatter’s Rights”: At any time, current French House resident members take precedence over non-residents, for as long as they have guaranteed housing through the housing office.

4Typically, however, the housing office gives the Language Houses more freedom than other dorms.

That is to say, all such resident members are promised continued residence in French House should they want it.^{5}

2. During the “summer” time, five spaces (or as many as are physically available) must be reserved for freshmen.^{6} These reserved spaces take precedence over all other non-residents who wish to live in French House! During “non-summer” times, these reservations do not apply.

This rule assumes that 5-506 is a triple, and that 5-507 is a double. If 5-507 should crowd, the new sixth space will be given to an incoming freshman. If upperclassmen wish to live in these rooms, an equal number of singles will be left open.

3. For any remaining spaces, at any time, preference is next given to former resident members of French House who have temporarily left the house (usually by leaving MIT, e.g. Junior Year Abroad programs, time spent at another college). This preference is only given to those returning members who have guaranteed housing. Multiple members who wish to return are ranked first by seniority (determined by the number of terms of guaranteed housing left out of eight), and then by random lottery.

4. During “non-summer” times, preference is then given to former resident members of French House who do not have guaranteed housing. Multiple members are ranked in the same manner as former resident members with guaranteed housing (seniority, then lottery).

5. At any time, preference for any remaining spaces is next given to social members who wish to move in to French House, in the following order:

(a) Class Size This is to balance the classes. Accept social members in the same class as a class which does not have five members (until that class has five members). If two classes have less than five members, preference is based on seniority. If there are more social members in a class than needed to reach five, preference is based on the further criteria below.

(b) Duration of Social Membership The duration of social membership is counted from the meeting at which the student became a social member. Social members admitted at the same meeting are lotteried for spots if there aren’t enough spots for all of them.

(c) Seniority Again, seniority is determined by the number of terms of housing left out of eight.

(d) Lottery

6. During the “summer,” preference is next given to other freshmen (beyond the five-freshman minimum above) who wish to live in French House. In this case, actual acceptance as a house member is determined by the Rush team.

7. At all times, if there are spaces remaining, the house should try to fill them with students who are sincerely interested in French House (usually friends of house members). Otherwise, MIT housing will temporarily fill the space with a student who will probably not be interested in French House.

5There is more on the subject of “squatter’s rights” in Section 9.2 below.

6Of course, current French House residents cannot be forced to move out to make room for freshmen!

8. During “non-summer” times, should any spaces remain, the house may (but is not required to) accept other house-to-house transfers. A word of caution: if no one in the house has heard of the person who wishes to move in, it is probably not a wise idea to allow this person in. The Vice-Président should instead try to convince the student that it would be better to become a social member first.

9.2 Placement in French House

NOTE: This section contains rules only for determining the order in which rooms are chosen, not who is admitted into French House.

9.2.1 Lottery List

The lottery list determines the order of room choice for people who wish to live in French House.

Spring lottery Sometime in April, usually in the week following officer elections, the Vice-Président makes the spring lottery list. This must be done in the presence of the Président and at least one other officer. All people eligible to live in the house, whether they actually will or not, are given a ranking in the spring lottery. People are randomly ordered within their class. Incoming seniors are given first priority, followed by juniors, then sophomores. Ninth and tenth term “superseniors” are lotteried with the incoming seniors. Social members and returning members are lotteried with their class. Non-members are placed at the bottom of the incoming sophomore class. Incoming freshmen are, of course, not in the spring lottery. Notwithstanding the above, a member’s priority in the lottery may be decreased as specified by duty implementation policies.

Rush lottery After the members of the incoming freshman class have been determined, the Ministre de R/O randomly lotteries them, ranking them after the incoming sophomore class. Note that although non-members may thus be ranked ahead of incoming freshmen in the lottery, the freshmen may take precedence for admittance into French House.

Lottery removal and insertion Anyone may request to be removed from the lottery list. This is not to be done carelessly! If, during the “summer,” this person wants or needs to be housed in French House, he is still to be considered a resident member (for purposes of “getting in” to the house), and is to be inserted back into the lottery list at the bottom of his class.

People who wish to be inserted into the lottery list after the Vice-Président has made the original list, are ranked at the bottom of the class in which they would have been placed in the spring lottery. Multiple such people within a class are ranked according to the time at which they made the request.

9.2.2 Choosing Rooms

After generating the spring lottery list, the Vice-Président prepares the list of people who will choose rooms, based on the number of available rooms in the house and Section 9.1, and ordered according to the lottery. Then the Vice-Président asks each person on this “room choice list,” in order, to choose a room.^{7} The Vice-Président has the authority to establish a reasonable limit to the amount of time a person may spend choosing a room, such that the room choices are completed by the deadline suggested by the housing office.

7This procedure doesn’t have to be door-to-door. He can post the lottery and a list of rooms, and request that people come to him with their room choices.

• Freshmen choose rooms at the end of Rush, after the Rush lottery. Co-ed rooms are prohibited, but otherwise the lottery ranking should be used. Most often, the freshmen just work out a compromise. All freshmen not living in crowded rooms are added to the room choice list.^{8}

• Anyone in the lottery who is anticipating a move out of French House before the end of “summer” should not be asked to choose a room, but will retain his ranking in the lottery. This person should be nearly certain that he will not be living in French House, since otherwise it will be necessary to partially redo room choices.

• If a room is vacated (e.g. someone moves out or graduates early) and there are crowded rooms in French House, these rooms de-crowd. Room 5-507 will de-crowd first, then 5-506 (both to doubles). People within a room are asked to de-crowd based on lottery ranking, unless they work out something else themselves. If nobody wishes to de-crowd, the housing office will probably take away the lower housing rates. The person who left the crowded room will be inserted into the room choice list. After any de-crowding, if there is still a vacant room, the Vice-Président will insert the non-resident who, by Section 9.1, is “next in line” to live in French House, into the room choice list.

Next, from the new room choice list, beginning with either the newly inserted person or the person ranked just below the person who vacated the room (whoever is ranked higher), the Vice-Président will give members the option of moving to the vacant room. Note that “squatter’s rights” prohibit members from choosing rooms that are not vacant.

10 House Meetings

House meetings are called and run by the Président (or the Vice-Président in the former’s absence). Meetings are usually held once per month. There is generally a predefined agenda, which the Président should post along with meeting announcements at least a week before the meeting date. Everyone is expected to attend the entire meeting unless prior commitments do not allow it.

Should a house member wish a meeting, he should talk to the Président. If it is a valid request, the Président will call a meeting.

11 Voting Policy

The following policy describes the methods and types of voting which must be observed during house meetings and at other times when it is needed (such as balloting or poll votes).

Every member of French House has the right to vote.

11.1 Terminology

A simple majority is defined to be more than one-half of a group. A two-thirds majority is defined to be 2/3 or more of a group. The presiding officer is the officer who conducts the house meeting.

8Both 5-506 and 5-507 are considered crowded as triples, according to the housing office.

11.2 Quorum

Any vote taken at a house meeting is not binding unless quorum exists at the time of the vote. The requirement of quorum cannot be suspended.

Quorum exists if a majority of the house members are present at the time of the vote.

11.3 Voting Procedures

11.3.1 General Procedures

Except for the cases enumerated below, the presiding officer decides the method and type of voting for every individual vote. If there is an objection to either the method or type of voting, a simple majority vote by show of hands is required to decide which method and type of voting to use.

After establishing the method and type of voting, the presiding officer declares the issue being voted upon, enumerating all voting options. The vote is then taken, and the presiding officer announces the official results. If the vote is not by secret ballot, a house member has the opportunity to change his vote any time before the official results are announced.

11.3.2 Specific Voting Procedures

Issues specifically requiring a two-thirds majority vote are considered fairly important. Therefore, the Président must announce all issues requiring a two-thirds majority vote at least one week before the vote is taken. If an unexpected or urgent issue arises requiring a two-thirds majority vote, and is not announced in time, then the issue may be voted upon only if the house first votes by simple majority to allow the two-thirds majority vote.

• Officer elections require simple majority votes by secret ballot. For more details, see Section 12.1.

• A vote on social membership requires a two-thirds majority vote by secret ballot.

• The decision of whether to make a “major” house purchase requires a two-thirds majority vote. The interpretation of the word “major” is left to the Président.

• Significant changes to the house guidelines and policies requires a two-thirds majority vote. The interpretation of the word “significant” is left to the Président.

11.4 Methods of Voting

11.4.1 Show of Hands

This is the most common method of voting, in which votes for a particular choice are indicated by a raised hand. The presiding officer enumerates all choices (ending with a call for abstentions), and pauses after each while he and the secretary both count raised hands.

11.4.2 Secret Ballot

In this method of voting, each member is given a slip of paper on which to write his vote. The final results must be tallied by the presiding officer and at least one other officer. Blank slips of paper, illegibly written votes, and votes for options other than those under consideration, are counted as abstentions.

Secret ballot votes, once cast, cannot be changed.

11.4.3 General Consent / “White Ballot”

General consent voting is used when it is thought that members are more or less in agreement on an issue. After calling for a “white ballot” vote and announcing the motion under consideration, the presiding officer asks for any objections to the method of voting. If there are no objections, the motion passes. If a member objects, discussion will continue, and general consent voting cannot be attempted again on the same motion.

A point of information does not constitute an objection, unless it results in further discussion.

11.4.4 Informal “Straw” Poll

This is a non-binding, unofficial vote which anyone may request (at the discretion of the presiding officer) during discussion to see how members stand on the issue under discussion. Quorum is not required for this method of voting.

11.5 Types of Votes

11.5.1 Plurality Vote

In a plurality vote, a voter is asked to choose among more than two options (or abstain). The option garnering the greatest number of votes passes.

If the vote is close, or the number of votes cast is small, the presiding officer may decide to repeat the vote, disregarding options which received few votes. The presiding officer may also decide to switch to another type of voting.

Plurality votes are generally used for minor issues like choosing a day on which to hold a social event.

11.5.2 Simple Majority Vote

In a simple majority vote, one choice must receive a simple majority of the votes cast in order to pass. A tie vote between two options, if not broken by the presiding officer (see Section 11.7), results in neither option passing. The only exception to this rule is a simple “yes/no” vote on a motion, in which case a tie means that the motion fails.

If there are more than two options among which to choose, and no option receives a majority of the vote, then the option receiving the fewest number of votes is removed from consideration, and the remaining options are voted upon again.^{9} The number of remaining options must always be two or greater; that is, an option cannot pass simply because all others were eliminated. If the number of options cannot be reduced to two, the presiding officer will decide whether to call the vote undetermined, to repeat the most recent vote, or to begin again with all options and repeat the process of elimination.

Simple majority votes are used for most house decisions.

11.5.3 Two-Thirds Majority Vote

In a two-thirds majority vote, one choice must receive a two-thirds majority of the votes cast in order to pass. If there are two options and neither receives the necessary two-thirds majority, neither option passes. The only exception to this rule is a simple “yes/no” vote on a motion, in which case the motion fails unless the “yes” votes constitute a two-thirds majority.

9If multiple options tie for fewest number of votes, they are all eliminated.

If there are more than two options among which to choose, and no option receives the necessary majority, the process of elimination described in Section 11.5.2 is used.

Two-thirds majority votes are used for more important house decisions. See Section 11.3.2 for examples.

11.6 Abstentions

It is every member’s right to abstain during a vote. An abstention is not to be considered a “nay” vote, but rather as a vote not cast. That is, the total number of votes cast does not include abstentions.

A voter should abstain if he knows nothing of the issue being voted upon (for instance, if the person arrives just before the vote is taken), or if the vote involves direct personal or financial gain.

11.7 Vote of the Presiding Officer

In general, the presiding officer does not vote, although his presence counts toward quorum. The presiding officer may, however, vote when the voting is sufficiently close that an extra vote might alter the outcome. For example, if during a simple majority vote there are 50 “yes” votes and 49 “no” votes, the presiding officer is allowed to vote, for another “no” vote would result in a tie and thus a failed motion. Although technically the presiding officer may use this privilege to purposely tie a vote, the spirit of the rule is to avoid deadlock by permitting a tie vote to be broken.

The presiding officer is always allowed to vote during secret ballot voting, and in this case cannot function as a tiebreaker.

11.8 Proxies

A proxy is a house member who is given the power to vote for someone who cannot attend the meeting. There are no restrictions on the proxy’s use of this extra vote. The presiding officer can neither designate a proxy nor be a proxy.

A member who wishes to designate a proxy must seek approval from the Président before the beginning of the meeting, explaining the inability to attend the meeting and naming the proxy. The Président should not approve the request if he feels that the reason for the member’s absence does not merit the privilege of having a proxy.

For the purpose of determining quorum, the person who is a proxy is still counted only once.

11.9 Absentee Ballots

An absentee ballot is a vote on a specific issue which has been announced prior to the meeting. The vote pertains only to one specific issue.

A member who wishes to cast an absentee ballot must seek approval from the Président before the beginning of the meeting, explaining the reason for his absence and giving the Président a piece of paper with a description of the issue, the member’s vote, and his signature. The Président should not approve the request if he feels that the reason for the member’s absence does not merit the privilege of casting an absentee ballot.

During the course of the meeting, the nature of the issue under discussion may change. The presiding officer will determine if the issue has changed so much that the absentee ballot is no longer relevant. If this is the case, the absentee ballot will count as an abstention.

Absentee ballots do not count towards quorum.

12 Officers

12.1 Elections

Elections for house officers are carried out at the first house meeting in April. At least one week before the meeting, the Président will post nomination sheets. House members may nominate themselves for any position. Nominations for multiple offices are permissible. Nominations for an office may also be made during the meeting and before the actual election for that office. The only requirement is that officers must reside in French House for the academic year following their election.

The order of the vote follows the order in which the officers are listed below. The voting procedure is outlined in Section 11. After an office election, the losers have the opportunity to be nominated for further offices.

The Président, Vice-Président, Trésorier, and Ministre des Aliments must all be separate individuals.

It is possible for people to run together for an office. For officers who are exempted from other house duties (e.g. the Ministre des Aliments and the Trésorier), the pair is expected to perform the same duties as one officer and one non-officer. For example, two Ministres des Aliments would cook once a week (the same day) between them and jointly carry out the duties of their office.

Usually two Ministres de R/O are elected^{10}. It is advisable that the Ministres de R/O run as a team, since they must work well with each other. The number of Représentants à la Maison Neuve is determined by the New House Consitution.

After elections, all offices are supposed to have immediate takeover. The Trésorier often does not take over until the following year, for simplicity in end-of-year bookkeeping. All offices should have an apprenticeship period.

12.2 Président (President)

• Organize, announce, and conduct house meetings

• Act as a liaison between French House and the MIT administration

• Represent French House in all external matters, when someone else hasn’t specifically been given such responsibility (e.g. the Vice-Président, the Ministre des Aliments, etc.)

• Receive all mail addressed to French House and direct it to the appropriate officers (inform the dorm mail stuffers)

• Mediate house and personal disputes as impartially as possible

• Maintain tâche lists, organize nettoyage schedules, dispatch end-of-year cleanup duties, and take appropriate punitive actions if necessary to ensure fulfilment of individual duties

• Maintain the lmf locker and the mailing lists

• Work with the Secrétaire to inform new French House members of the rules

• Work with all other officers to coordinate their responsibilities and activities, and ensure that the other officers are fulfilling their duties

• After elections, organize the senior brunch

10They are also usually both freshmen/returning sophomores.

12.3 Vice-Président / Ministre de Logement (Vice-President / Housing Chair)

• Assume any responsibilities of the Président when he cannot fulfill them, at the Président’s request

• Represent French House in all MIT housing matters

• Handle all internal housing affairs, including conducting the spring housing lottery

• Keep the Président informed of housing and any other pertinent matters

• Record non-monetary penalties specified by duty implementation policies

• Is often here for Rush

12.4 Trésorier (Treasurer)

• Do signout sheets about every two weeks

• Do the dîner chart every week

• Credit personal accounts from receipts

• Pay IM fines

• Pay meat, vegetable, milk, and other ordered-food bills

• Order Technique photos and house shirts

• Pay for Technique photo when it’s taken

• Keep track of receipts, invoices, bank statements, and other money records

• Keep track of money from International Fair

• Make deposits

• Receive Institute house tax return (from New House) - not credited to non-resident French House members

• Assess monetary penalties specified by duty implementation policies

• Calculate for house bills:

– dinner costs - charged per meal

– social costs - charged to all

– pause-cafés - can be changed to everyone or per item (you eat, you pay)

– Rush costs - not charged to incoming freshmen

– senior brunch costs - not charged to seniors

– house costs - charged to all

• Make people pay!

• NOTE: Social members are usually billed less than resident members (usually around 75% of resident member bill), but they do not get the house tax refund

• NOTE: the Trésorier is exempt from cooking duties

12.5 Ministre des Aliments (Food Steward)

• Organize and manage cooking teams

• Coordinate activities related to cooking (e.g. distribution of and delivery of food, nettoyage)

• Purchase items for signout

• Purchase items for menus

• Start up and maintain accounts with food services

• Designate items as “house” or “cooking,” with house consent

• Designate days for cooking and not cooking, with house consent

• Bring up any issues dealing with house money to house meetings

• Must be consulted in all matters concerning outside purchases of goods necessary for house functions

• Must limit purchases to house activities and other activities at the joint discretion of the Ministre des Aliments and the Trésorier

• Must implement all guidelines approved at house meetings

• Must jointly agree with the Trésorier on all purchases exceeding $20 per single unit item • NOTE: the Ministre des Aliments is exempt from cooking duties

12.6 Secrétaire

• Take minutes at all house meetings, or appoint a proxy if attendance is impossible

• Post the minutes from the previous house meeting

• Bring the minutes from at least the last seven to eight house meetings to every meeting for reference purposes

• Work with the Président to inform new French House members of the rules

12.7 Ministre des Sports (Athletic Chair)

• Manage the house intramural sports program

• Stop by the athletic office as often as possible to check for new sport fliers, schedules, fines, and any other pertinent information

• Post sports information for the entire house, espescially upcoming sports and sport schedules • Post signup sheets for each sport, and fill out rosters and choose captains and referees, if

needed, for sports that have enough people signed up

• Attend the monthly athletic chair meetings (lest the house be fined!)

• Pay any fines (incurred by missed meetings or games) promptly either by paying for them personally and having one’s house bill reduced, or by getting a check for the fine from the Trésorier

12.8 Ministre de R/O (Rush Chair)

• Talk to New House Rush Chair to find out how many early returns French House has

• Determine who will get the early returns. The Ministre de R/O asks who wants to come back to work Rush and tries to get as many people to come back as possible. Additional people can usually return by rooming with early returns or summer residents.

• Meet with New House and Language House rush chairs to coordinate Rush activities

• Organize a summer Rush mailing

• Order and buy food, put up posters, etc. The Ministre des Aliments should be asked about the possibility of ordering food since it is more convenient.

• Plan, organize, and run French House Rush. Anyone who gets an early return must work all of French House Rush.

• Have a meeting with Rush workers (those who help significantly with Rush) and decide which freshmen (seem interested / could survive in the house / speak enough French / seem interested in learning French) enough to live in the house. The meeting should be after the mandatory brunch.

12.9 Ministre des Soirées (Social Chair)

• Organize social events, most commonly parties, at least one per month, including the Tower of Babel party and a semi-formal once a year

– Pick a location, time and date. Usually the fifth floor lounge or the first floor of House 6 (if it is okay with the Germans), Friday or Saturday, 9 p.m. to 1 a.m. (check for conflicts with other activities that many house members participate in)

– Advertise. Usually posters, word of mouth, and e-mail (at least 2 to 3 days in advance) are used. Also be sure to post it to the nhsocial mailing list so that other New House social chairs can take it into account.

– Get food and drinks. Put a sign-up sheet up to get help with the food. It is suggested to keep a reserve of Kool-Aid in case drinks run out.

– Arrange for the music. Post a sign-up sheet for DJs. A good rule of thumb is to have a new DJ every hour or so. Borrow CDs from people in the house before the party. Borrow the stereo equipment; try to do this early.

– Set up and decorate.

– Make sure someone cleans up after the party, so it looks the same as (or better than) when you arrived.

• May organize trips to Talbot House, Québec, or similar excursions

• Inform the house of other known social activities, especially New House events

12.10 Ministre de l’Education (Education Chair)

• Actively encourage the use of French in the house through activities such as Mot du Jour and the French marathon

• Inform the house of various French-related educational, cultural, and social events in the area

12.11 Ministre de l’Environnement (Environment Chair)

• Represent French House in matters of New House recycling

• Coordinate the house’s recycling activities

• Maintain the house plants, posters, and other decorations

• Take down posters if house is not open during the summer

12.12 Représentant à la Maison Neuve (Representative to New House)

• Attend New House meetings, or secure a proxy

• Vote as you believe represents the house, not just your own personal views

• Tell house members what happens at the meetings (through postings or meetings)

• Get people’s input and ideas and bring them up at the meetings

• Try to get people involved and enthusiastic

13 Traditions

• The Oar “The oar” (“Notre Râme”) is shared between French House and Spanish House because it was broken by a joint Spanish House-French House IM Crew boat. There are well-defined rules about sharing it. One house may remove it from the other only on Friday, before midnight. The oar may only be moved once each Friday. A house may not prevent the other from taking it at the right time. This includes hiding it.

• The French House Dance Usually happens at least once a party. Danced to “The Love Shack.”^{11}

• The French Marathon Every year there is a French Marathon, organized by the Ministre d’Education. During the marathon, everyone speaks nothing but French. The exact rules are set when it is done, but you are not required to speak French when inside someone’s room.

• Tower of Babel Party Not really a French House tradition but one of the Language Houses. Each year (since 1977), the Language Houses have a party (or another event) together. Sometimes an actual tower is built.

11Epsilon Theta has a similar dance for “Beat It.” Our dance is rumored to have been imported from Alpha Epsilon Pi.

• Dîner de Noël Near the end of the first term, the Dîner de Noël is held. This is a big dinner to which house alumni are invited. People dress nicely.

• Secret Santa (Père Noël) For two weeks before the Dîner de Noël, house members can be Secret Santas. Everyone who is participating draws the name of another and over the two weeks gives their “Santee” notes, small gifts, etc. For the Dîner de Noël, a more substantial gift is wrapped up and the Secret Santas are revealed at the end of dîner (after the Santee has had 3 guesses).

• Mystery Night This is a recent tradition. A few members of the house organize a role-playing murder mystery which takes place in the house over the course of a few hours. Everyone who wants to participate is given a character to play, Assassin’s Guild style.

• International Fair Usually French House has a entry in the International Fair, an event which takes place toward the end of the second term. We sell croissants and other French foods out on Kresge Oval.

• Semi-Formal French House holds a semi-formal dance party in the spring. It gives people a chance to dress up more than usual. Dates are not necessary, and it’s nice to specially invite friends. It is normally not held in French House; in past years it has been held in Tang Hall and in McCormick, depending on where we have connections.

• Senior Brunch Near the end of the year, the seniors are treated to the senior brunch. The newly-elected president organizes it. The seniors dress nicely, everybody should attend, and after the brunch, a modest gift is given to each senior.